OpenStudy (ress):

What values of x and y satisfy −13/3x−2y=7 and 3x+2/5y=5 simultaneously

OpenStudy (triciaal):

one approach multiply by the LCD to get rid of the fractions put each in the form y = set the y's equal and solve for x substitute this value in the expression to compute the value of y

OpenStudy (queenie):

-13/3x - 2y = 7...multiply by 3 to get rid of fractions 3x + 2/5y = 5...multiply by 5 to get rid of fractions -13x - 6y = 21 15x + 2y = 25...multiply by 3 ------------- -13x - 6y = 21 45x - 6y = 75 (result of multiplying by 3) ------------add 32x = 96 x = 96/32 x = 3 15x + 2y = 25 15(3) + 2y = 25 45 + 2y = 25 2y = 25 - 45 2y = - 20 y = -10 solution is : (3,-10)